The personnel manager of a company feels that 60% of the employees will have enhanced
skill after attending the training programme if they are sponsored for an in-house training
programme. A sample of records of 49 employees of the company, who attended the
training programme on skill development, reveals that only 24 of them are having enhanced
skill after attending the training programme. Find the probability that the sample of
employees who attended the training programme have enhanced their skill.

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

point estimate mss 24/49=0.4898
z=(0.4898-0.60)/sqrt(0.6*0.4/49)
=-0.1102/0.07
=-1.57
that probability that z is <= -1.574=0.0577
(remember this is a continuous function, so the probability has to be over an interval, not a single value on the curve, where the probability would be 0.)

Answer by ikleyn(52778)

.

How this problem is worded, it is a soup of words.